==================================================
Yves Perrin
+ Actions done and progress
- helped S. Lueders/IT-CO to configure the trackers of his savannah
project
- iterated with Mathias Philippe Dutour who wants to extract data from
Savannah and correlate it automatically to the CVS tags of the software
components --> he will use the export facility. I will enhance it for
daily execution
- wrote a parser to extract and display the NB summary from its XML
representation.
- working on its integration in the SPI twiki with automatic statistics
display on 'mouseover'
- created many new forums for Atlas HN.
- continue putting pressure on people to fix wrong email addresses in
savannah (generating rejected mail messages)
- helped a savannah user who messed up with his login and AFS
authentication
- deleted a set of SPAM comments which managed to get into Savannah
- continue putting pressure on PH authors to follow CERN conventions
when registering new web sites (names, categories, etc)
- involved in discussions with Stephan Haywood concerning the HN support.
I will continue with the day to day administration, possibly writing
a few utility script to help with the administration but I can't help
with the technical support of HN (lack of time, of knowledge of the
HN inside, and of organization for source control).
+ Next actions
- complete the integration of the NB summary in the SPI twiki (as well
as the few remaining SPI web pages)
+ AOB
- will be away Monday 6 August
==================================================
Bertrand Bellenot
+ Actions done
- rebuild globus plugins for ROOT v5.14.00g on
slc4_ia32_gcc34 and slc3_ia32_gcc323. There
are still some problems on amd64, due to the
weird way globus distributes its binaries...
+ Next actions
- try to build globus plugins on slc4_amd64_gcc34
as soon as a solution has been found...
==================================================
Piotr Kolet
+ Actions done
- Nightlies builds
--all updates are made in scripts-devel directory, then they are commited and updated in script directory, that are one in production.
-- nighlies on Windows - fixes,
-- Mac OSX on intel. Preparing the platform, preparing the web page, fixing mac2afs script, etc.
- trying to solve the COOL test problem with Marco
==================================================
Stefan Roiser
+ Actions done
- Hypernews moved successfully, All LHC instances are now located
on lcgapp02 (lcgapp03 being hot standby).
- problem with sync to fallback host (ssh in cron not working)
- External s/w installations for mac/intel
gsl 1.8
- Interactions with CORAL and POOL to remove errors/warnings for mac/intel builds
- Several interactions with LCG developers to cure problems in the nightly builds
+ Next actions
- External s/w
gccxml 0.6.0_patch4
qt 3.3.8
pyqt 3.17.2
+ Aob
- osx104_ppc_gcc401(_dbg) and slc3_ia32_gcc323_dbg have been abandoned as LCG platform
- osx104_ia32_gcc401(_dbg) are now official LCG platforms
- for LCG installations a symlink slc4_XXX_gcc34(_dbg) -> slc4_XXX_gcc41(_dbg) shall be provided
There are minutes attached to this event.
Show them.